# Moin et al

### v2

@@ Abstract
[Abstract]Various architecture frameworks for software, systems, and enterprises have been proposed in the literature. They identified several stakeholders and defined modeling perspectives, architecture viewpoints, and views to frame and address stakeholder concerns. However, the stakeholders with data science and Machine Learning (ML) related concerns, such as data scientists and data engineers, are yet to be included in existing architecture frameworks. Only this way can we envision a holistic system architecture description of an ML-enabled system. Note that the ML component behavior and functionalities are special and should be distinguished from traditional software system behavior and functionalities. The main reason is that the actual functionality should be inferred from data instead of being specified at design time. Additionally, the structural models of ML components, such as ML model architectures, are typically specified using different notations and formalisms from what the Software Engineering (SE) community uses for software structural models. Yet, these two aspects, namely ML and non-ML, are becoming so intertwined that it necessitates an extension of software architecture frameworks and modeling practices toward supporting ML-enabled system architectures. In this paper, we address this gap through an empirical study using an online survey instrument. We surveyed 61 subject matter experts from over 25 organizations in 10 countries.
[…]
keywords:
@@ 1 Introduction
Architecture frameworks provide conventions, principles, and practices for architecture descriptions in a particular application domain or stakeholder community 1. There exist several well-established examples, including The Open Group Architecture Framework (TOGAF) 2, 3, the U.S. Department of Defense Architecture Framework (DoDAF) 4, the Treasury Enterprise Architecture Framework (TEAF) 5, the British Ministry of Defence Architecture Framework (MODAF) 6, the Zachman Framework 7, the “4+1” View Model of Software Architecture 8 and its updated version with the Decision View 9, as well as the Reference Model of Open Distributed Processing (RM-ODP) 10, 11, 12, 13. TOGAF, DoDAF, TEAF, and MODAF were primarily concerned with enterprise architectures. MODAF was replaced by the NATO Architecture Framework (NAF), which in its fourth version (NAFv4) 14 provided guidance not only on describing enterprise architectures but also system architectures for military and business use. Furthermore, the Zachman Framework, a generic framework for information systems and enterprise architectures, and the “4+1” View Model of Software Architecture were among the early works that offered the foundations for the more recent architecture frameworks. Last but not least, RM-ODP was not merely a reference model but a set of four international standards, including an architecture framework for distributed information processing in heterogeneous environments. Each of the above-mentioned architecture frameworks identified stakeholders, such as end-users, software developers, system integrators, system engineers, business domain experts, executives, and other corporate functions 8, 3. Also, they defined architecture viewpoints and views. Section 2 provides some background on architectural artifacts, such as viewpoints and views.
@@
However, prior work in architecture frameworks did not recognize any stakeholder with ML-related concerns, although ML is increasingly making software and information systems smart, and ML components are assuming a prominent role in many systems and organizations. Consequently, existing architecture frameworks lack any viewpoint or view dedicated to the stakeholders of ML artifacts. Hence, we argue that architecture descriptions of smart (i.e., ML-enabled) software systems, which need to reflect on their ML components and the interactions of the ML components with the non-ML components, cannot be adequately designed using state-of-the-art architecture frameworks. Essentially, ML is a separate field, a sub-discipline of Artificial Intelligence (AI) rather than Software Engineering (SE). Thus, it has its vocabulary, skill set, and know-how, which are different from the ones possessed by typical software developers. Recent research work, such as the interviews conducted by Nahar et al. 15, stressed the necessity of collaboration between software engineers and other specialists, such as data scientists, for building ML-enabled systems and the associated challenges. In particular, they emphasized the human factors of collaboration, including the need to separate the data science and SE work, as well as to coordinate between them, negotiate and document interfaces and responsibilities, and plan the system operation and evolution. According to them, those human collaboration challenges appeared to be the primary obstacles in developing ML-enabled systems. Additionally, past work mainly focused on ML models, such as the challenges of learning, testing, or serving ML models. They were rarely concerned with the entire system, with many non-ML parts into which the ML model is embedded as a component, which requires coordinating and integrating work from multiple experts or teams 15. Similarly, Lewis et al. 16 elaborated on the so-called ML Mismatch problem that typically occurs in the process of development, integration, deployment, and operation of ML-enabled systems due to incorrect assumptions made about system elements by different stakeholders, such as data scientists, software engineers, and the operations team.
@@
In this paper, we postulate that ML artifacts of smart systems deserve to be treated as having a separate identity, distinguished from software source code and raw data. Moreover, stakeholders should see ML aspects of systems in different ways (i.e., using tailored notations and at various detail levels) such that they can better understand architecture descriptions of ML-enabled systems. In this way, they should become capable of efficient communication with other stakeholders and collaborate to contribute to the system or its architecture description, for example, by rigorous requirement elicitation and tracing. For instance, a data scientist is often in charge of analytics modeling and is primarily concerned with ML performance metrics, such as Accuracy, Precision, and Recall of the ML model when faced with unseen test data. However, a data engineer is responsible for analytics operations, thus ensuring scalable data processing 17. By contrast, other stakeholders, such as software engineers, software architects, database engineers, and system engineers, consider other aspects of performance that might be affected by the performance of the ML component, for example, by the delay introduced as a result of the predictions of the ML model, or might be unrelated to the ML component. The same holds for security: software and system engineers might underestimate the potential vulnerability of ML-enabled systems through adversarial attacks on the ML models. Therefore, data scientists and software engineers often have different notions of security. There also exists other challenges, such as the need for versioning ML-artifacts, for example, ML models, or at least their parameters (in the case of parametric ML models) and hyperparameters, monitorability concerns for ML models, as well as new Ethical, Legal, and Social Implications (ELSI), for example, new privacy challenges, or the explainability, fairness, and trustworthiness of ML components.
@@
We argue that existing software system architecture perspectives and viewpoints cannot realize the above-mentioned requirements. As opposed to classic deterministic or stochastic behavioral, functional, or logical models of software systems, ML models rely on inference on data. Apart from certain application scenarios, in most cases, ML models will need to be retrained on new data in the future or even continuously trained in an online learning scenario. For instance, they can become fooled, biased, or compromised through adversarial attacks carried out on the training data that are fed to them. While existing perspectives and viewpoints may provide model kinds and architecture views that enable data flow modeling, they essentially lack a systematic approach to inference models and mechanisms, such as mathematical formalisms that can capture the underlying statistical models or Data-Flow Graphs (DFGs), also known as Computational Graphs (CGs) 18. Moreover, without using the model kinds, formalisms, and notations that are commonly used by ML experts and data scientists, architecture models cannot fully support a fruitful stakeholder discussion and communication, nor can they be used for system documentation or automated code generation. Those are key use cases for system architecture models.
@@
The remainder of this paper is structured as follows: Section 2 provides a brief background, whereas Section 3 reviews the related work in the literature. Further, Section 4 elaborates on the research design and methodology. Moreover, Section 5 proposes new stakeholders, viewpoints, and views. Also, Section 6 points out possible threats to validity. Finally, Section 7 concludes and suggests future work.
2 Background
The ISO/IEC/IEEE 42010:2011 standard for architecture descriptions in systems and software engineering 1 defines the architecture of a system as fundamental concepts or properties of the ‘system in its environment embodied in its elements, relationships, and in the principles of its design and evolution’ 1. An Architecture Description (AD) is a work product that expresses an architecture. Every system stakeholder has various concerns regarding the System under Consideration (SuC), also known as the System of Interest (SoI), in relation to its environment. Concerns arise throughout the life cycle of the system from the system requirements, design choices, implementation, and operations. Performance, reliability, security, privacy, distribution, openness, evolvability, modularity, cost, and regulatory compliance are a number of examples of concerns 1, 3. Furthermore, Separation of Concerns (SoC), which is a vital design principle in SE, can be applied at different abstraction levels. At a lower level, it is interpreted as the modularity of the software system implementation, with information hiding and encapsulation in modules that have well-defined interfaces. However, at a higher level, it means describing the system architecture from the perspective of different sets of concerns (i.e., different architecture viewpoints). This is what an architecture view does. It is a work product that expresses ‘the architecture of a system from the perspective of specific system concerns’ 1. Additionally, an architecture viewpoint is a work product that establishes ‘the conventions for the construction, interpretation, and use of architecture views to frame specific system concerns’ 1. In other words, a view is what a stakeholder can see, whereas a viewpoint is where the stakeholder is looking from (i.e., the perspective or template that determines what they should see). A viewpoint is generic and can be stored in a library for reuse. However, a view is specific to the architecture for which it is created 3. For instance, Table 1 and Figure 1 illustrate a sample architecture viewpoint and its corresponding architecture view, respectively, based on the TOGAF architecture framework. The view describes the architecture of a Cyber-Physical System (CPS) concerning the physical location of the sensors in the distributed system.
Table 1: A sample architecture viewpoint in TOGAF 3
@@ 2 Background
Figure 1: A sample architecture view in TOGAF 3 governed by the viewpoint of Table 1
@@
The “4+1” View Model of Software Architecture 8, which was proposed in 1995 and updated in 2009 9, considers the following architecture viewpoints [^1] Note that they used the term view in their work, which corresponded to the notion of viewpoint as defined by ISO/IEC/IEEE 42010:2011 1. for software systems: i) The logical viewpoint concentrates on the functionality of the software system. The key stakeholder group whose concerns are framed by this viewpoint is the end-user group. Moreover, the components in the views associated with this viewpoint are often classes. ii) The process viewpoint focuses on the non-functional aspects, such as performance, availability, fault tolerance, integrity, and scalability. This viewpoint is considered primarily for system integrators and system engineers. Furthermore, the components in the views associated with this viewpoint are typically tasks. iii) The development viewpoint is concerned with the software development and management aspects, such as the subsystems and modules’ organization, reuse, and portability. The main stakeholder groups here are the programmer group (i.e., software developers) as well as software managers. Also, the components in the views associated with this viewpoint are usually modules and subsystems. iv) The physical viewpoint deals with scalability, performance, and availability on the physical layer, thus considering the network communications, distribution, topology, and the mapping of software onto the hardware. The main stakeholder group here is the system engineer group. Moreover, the components in the views associated with this viewpoint are often nodes. v) The scenarios viewpoint ensures understandability and includes a small set of use case scenarios that can show how the elements of the above-mentioned viewpoints can work together. In fact, this viewpoint is an abstraction of the system requirements. It is worth mentioning that this viewpoint is in some sense redundant from the other ones. This is why the name contains ‘+1’. However, it serves the important purpose of being a driver to discover the architectural elements during the design phase and playing a validation and illustration role after the completion of the design for the test of an initial system prototype. This viewpoint is mainly intended for end-users and software developers. Lastly, the components of the views associated with this viewpoint are steps and scripts. vi) The decision viewpoint: Architects have various design choices and make a lot of design decisions. These decisions must be rigorously documented. The focus of this new viewpoint in the revised version of the original ‘4+1’ model is on design decisions.
@@
Further, the aforementioned standard 1 defined a model kind as a set of conventions for a type of modeling. An architecture viewpoint comprises one or more model kinds. Similarly, an architecture view comprises one or more architecture models. Additionally, each and every architecture viewpoint governs one or more architecture views; also, each and every model kind governs one or more architecture models. For instance, ‘data flow diagrams, class diagrams, Petri nets, balance sheets, organization charts, and state transition models’ are model kinds 1. Finally, architecting is the ‘process of conceiving, defining, expressing, documenting, communicating, certifying proper implementation of, maintaining, and improving an architecture throughout a system life cycle’ 1. It is typically conducted in the context of a project or an organization.
3 Related Work
@@
The “4+1” View Model of Software Architecture, as explained in Section 2, was an early but leading work concerned with the architecture of software systems. Other architecture frameworks, such as the Zachman Framework 7, which preceded this, or the other ones that succeeded it (see Section 1), had also identified a number of stakeholders, as well as several viewpoints and views. In the following, we refer to DoDAF 4 and TOGAF 2, 3 whose latest versions were published in 2010 and 2018, respectively, as two representative examples. In particular, we highlight their data-related aspects since those are the most related ones to the target of this study.
@@
DoDAF 4 defined several viewpoints including the Data and Information Viewpoint 1 (DIV-1, called conceptual data model), DIV-2 (called logical data model), and DIV-3 (called physical data model). DIV-1 was mainly intended for non-technical stakeholders and showed the high-level data concepts and their relationships. Relationships at this level would be simple (i.e., not attributed). However, DIV-2 bridged the gap between the conceptual and physical levels by introducing the attributes and structural business process (activity) rules that formed the data structure. This viewpoint was intended for system architects and analysts. For example, an Entity-Relationship (ER) diagram or a Class diagram could be deployed as a candidate model kind for this viewpoint. Lastly, DIV-3 provided the physical schema for the data and information and was close to the actual implementation level. This could often be represented by tables, records, and keys in a relational database management system or through objects in an object-oriented data model. The physical data model (i.e., DIV-3) was intended for database engineers, software developers, and system engineers.
@@
Moreover, the TOGAF standard 2, 3 is closely related to the Zachman Framework 7222In fact, there exists a mapping 2 between the TOGAF Architecture Development Method (ADM) and the Zachman Framework.. In its 2006 version (i.e., v8.1.1), TOGAF 2 adopted the data flow viewpoint [^3] Note that they used the term view in their work, which corresponded to the notion of viewpoint as defined by ISO/IEC/IEEE 42010:2011 1., which was intended for database engineers, ‘concerned with the storage, retrieval, processing, archiving, and security of data’, thus ‘assuring ubiquitous access to high-quality data’ 2. Further, the latest version of TOGAF (v9.2) from 2018 3 promoted a set of data-related viewpoints (which they called data diagrams): i) The conceptual data viewpoint had the key purpose of depicting the relationships between the critical data entities of an enterprise. This viewpoint was intended for business stakeholders. ER diagrams or simplified UML Class diagrams could be deployed as the model kinds supporting this viewpoint. ii) The logical data viewpoint was intended for application developers and database designers and showed the logical views of the relationships between the critical data entities within an enterprise. iii) The data dissemination viewpoint illustrated the relationship between data entity, business service, and application components. It showed how the logical entities were physically realized by application components. iv) The data security viewpoint intended to depict which actor (i.e., person, organization, or system) has access to which enterprise data. This could be demonstrated in a matrix form or as a mapping. Moreover, it could show compliance with data privacy laws and other applicable regulations. v) The data migration viewpoint showed the flow of data between the source and target applications when implementing a package or packaged service-based solution, for example, if a legacy application was to be replaced. Packages might have their own data model. Thus, data transformation might be necessary. The transformation could include data quality processes, namely data cleansing, matching, merging, and consolidating data from different sources, as well as source-to-target mappings. This viewpoint could be deployed for data auditing and establishing traceability. The detail level of the supporting views could vary. vi) The data life cycle viewpoint enabled managing business data throughout their life cycle from conception until disposal. Each change in the state of data should be shown through the views supporting this viewpoint. The data were considered entities in their own rights, decoupled from business processes and activities. This allowed common data requirements to be identified.
@@
Recently, various works have pointed out the need for upgrading architecture frameworks to support ML-enabled systems. Almost around the same time that we conducted our empirical study (2021), Muccini and Vaidhyanathan 19 published a related research paper based on their experience in architecting an ML-based software system. They identified the following stakeholders with ML-related concerns: data scientists, ML developers, ethics experts, and data engineers. Moreover, they envisioned the following concerns for them: ML model accuracy, model versioning, data quality, privacy, ethics, framework, and algorithmic choice. Our empirical results confirmed their identified stakeholders and had some overlap with their stated concerns.
Additionally, Lewis et al. 20 illustrated the role of architects in ML-enabled systems and highlighted some new challenges faced by architects and other stakeholders in devising software architecture of ML-enabled software systems, namely software systems that rely on one or more ML components (including ML models) to provide their capabilities. Their study was also published in 2021 when we conducted our survey. Their work emphasized the importance of bringing the SE and data science communities together to address the concerns of both stakeholder groups holistically. They also agreed with our understanding that ML artifacts must be treated as special software artifacts, and existing principles and practices in architecture frameworks must be updated to address new concerns emerging from these artifacts.
We started from our own experience to identify the gap in the literature and formulate the research questions. However, unlike most of the related work, we conducted an empirical study that referred to domain experts and asked them for their input, positions, and preferences on the matter. Note that Lewis et al. 16 also conducted practitioner interviews as part of their work. They showed that the development and deployment of ML-enabled systems involve three different perspectives, which include data science, software engineering, and operations. If they are misaligned due to incorrect assumptions, they cause ML mismatches, which can result in failed systems. Their study was also published in 2021 when we surveyed our experts.
4 Research design and methodology

### v3

# Enhancing Architecture Frameworks by Including Modern Stakeholders and their Views/Viewpoints

@@ Abstract
Various architecture frameworks for software, systems, and enterprises have been proposed in the literature. They identified several stakeholders and defined modeling perspectives, architecture viewpoints, and views to frame and address stakeholder concerns. However, the stakeholders with data science and Machine Learning (ML) related concerns, such as data scientists and data engineers, are yet to be included in existing architecture frameworks. Only this way can we envision a holistic system architecture description of an ML-enabled system. Note that the ML component behavior and functionalities are special and should be distinguished from traditional software system behavior and functionalities. The main reason is that the actual functionality should be inferred from data instead of being specified at design time. Additionally, the structural models of ML components, such as ML model architectures, are typically specified using different notations and formalisms from what the Software Engineering (SE) community uses for software structural models. Yet, these two aspects, namely ML and non-ML, are becoming so intertwined that it necessitates an extension of software architecture frameworks and modeling practices toward supporting ML-enabled system architectures. In this paper, we address this gap through an empirical study using an online survey instrument. We surveyed 61 subject matter experts from over 25 organizations in 10 countries.
Index Terms:
@@ I Introduction
Architecture frameworks provide conventions, principles, and practices for architecture descriptions in a particular application domain or stakeholder community [1]. There exist several well-established examples, including The Open Group Architecture Framework (TOGAF) [2, 3], the U.S. Department of Defense Architecture Framework (DoDAF) [4], the Treasury Enterprise Architecture Framework (TEAF) [5], the British Ministry of Defence Architecture Framework (MODAF) [6], the Zachman Framework [7], the “4+1” View Model of Software Architecture [8] and its updated version with the Decision View [9], as well as the Reference Model of Open Distributed Processing (RM-ODP) [10, 11, 12, 13]. TOGAF, DoDAF, TEAF, and MODAF were primarily concerned with enterprise architectures. MODAF was replaced by the NATO Architecture Framework (NAF), which in its fourth version (NAFv4) [14] provided guidance not only on describing enterprise architectures but also system architectures for military and business use. Furthermore, the Zachman Framework, a generic framework for information systems and enterprise architectures, and the “4+1” View Model of Software Architecture were among the early works that offered the foundations for the more recent architecture frameworks. Last but not least, RM-ODP was not merely a reference model but a set of four international standards, including an architecture framework for distributed information processing in heterogeneous environments. Each of the above-mentioned architecture frameworks identified stakeholders, such as end-users, software developers, system integrators, system engineers, business domain experts, executives, and other corporate functions [8, 3]. Also, they defined architecture viewpoints and views. Section II provides some background on architectural artifacts, such as viewpoints and views.
@@
However, prior work in architecture frameworks did not recognize any stakeholder with ML-related concerns, although ML is increasingly making software and information systems smart, and ML components are assuming a prominent role in many systems and organizations. Consequently, existing architecture frameworks lack any viewpoint or view dedicated to the stakeholders of ML artifacts. Hence, we argue that architecture descriptions of smart (i.e., ML-enabled) software systems, which need to reflect on their ML components and the interactions of the ML components with the non-ML components, cannot be adequately designed using state-of-the-art architecture frameworks. Essentially, ML is a separate field, a sub-discipline of Artificial Intelligence (AI) rather than Software Engineering (SE). Thus, it has its vocabulary, skill set, and know-how, which are different from the ones possessed by typical software developers. Recent research work, such as the interviews conducted by Nahar et al. [15], stressed the necessity of collaboration between software engineers and other specialists, such as data scientists, for building ML-enabled systems and the associated challenges. In particular, they emphasized the human factors of collaboration, including the need to separate the data science and SE work, as well as to coordinate between them, negotiate and document interfaces and responsibilities, and plan the system operation and evolution. According to them, those human collaboration challenges appeared to be the primary obstacles in developing ML-enabled systems. Additionally, past work mainly focused on ML models, such as the challenges of learning, testing, or serving ML models. They were rarely concerned with the entire system, with many non-ML parts into which the ML model is embedded as a component, which requires coordinating and integrating work from multiple experts or teams [15]. Similarly, Lewis et al. [16] elaborated on the so-called ML Mismatch problem that typically occurs in the process of development, integration, deployment, and operation of ML-enabled systems due to incorrect assumptions made about system elements by different stakeholders, such as data scientists, software engineers, and the operations team.
@@
In this paper, we postulate that ML artifacts of smart systems deserve to be treated as having a separate identity, distinguished from software source code and raw data. Moreover, stakeholders should see ML aspects of systems in different ways (i.e., using tailored notations and at various detail levels) such that they can better understand architecture descriptions of ML-enabled systems. In this way, they should become capable of efficient communication with other stakeholders and collaborate to contribute to the system or its architecture description, for example, by rigorous requirement elicitation and tracing. For instance, a data scientist is often in charge of analytics modeling and is primarily concerned with ML performance metrics, such as Accuracy, Precision, and Recall of the ML model when faced with unseen test data. However, a data engineer is responsible for analytics operations, thus ensuring scalable data processing [17]. By contrast, other stakeholders, such as software engineers, software architects, database engineers, and system engineers, consider other aspects of performance that might be affected by the performance of the ML component, for example, by the delay introduced as a result of the predictions of the ML model, or might be unrelated to the ML component. The same holds for security: software and system engineers might underestimate the potential vulnerability of ML-enabled systems through adversarial attacks on the ML models. Therefore, data scientists and software engineers often have different notions of security. There also exists other challenges, such as the need for versioning ML-artifacts, for example, ML models, or at least their parameters (in the case of parametric ML models) and hyperparameters, monitorability concerns for ML models, as well as new Ethical, Legal, and Social Implications (ELSI), for example, new privacy challenges, or the explainability, fairness, and trustworthiness of ML components.
@@
We argue that existing software system architecture perspectives and viewpoints cannot realize the above-mentioned requirements. As opposed to classic deterministic or stochastic behavioral, functional, or logical models of software systems, ML models rely on inference on data. Apart from certain application scenarios, in most cases, ML models will need to be retrained on new data in the future or even continuously trained in an online learning scenario. For instance, they can become fooled, biased, or compromised through adversarial attacks carried out on the training data that are fed to them. While existing perspectives and viewpoints may provide model kinds and architecture views that enable data flow modeling, they essentially lack a systematic approach to inference models and mechanisms, such as mathematical formalisms that can capture the underlying statistical models or Data-Flow Graphs (DFGs), also known as Computational Graphs (CGs) [18]. Moreover, without using the model kinds, formalisms, and notations that are commonly used by ML experts and data scientists, architecture models cannot fully support a fruitful stakeholder discussion and communication, nor can they be used for system documentation or automated code generation. Those are key use cases for system architecture models.
@@
The remainder of this paper is structured as follows: Section II provides a brief background, whereas Section III reviews the related work in the literature. Further, Section IV elaborates on the research design and methodology. Moreover, Section V proposes new stakeholders, viewpoints, and views. Also, Section VI points out possible threats to validity. Finally, Section VII concludes and suggests future work.
II Background
The ISO/IEC/IEEE 42010:2011 standard for architecture descriptions in systems and software engineering [1] defines the architecture of a system as fundamental concepts or properties of the ‘system in its environment embodied in its elements, relationships, and in the principles of its design and evolution’ [1]. An Architecture Description (AD) is a work product that expresses an architecture. Every system stakeholder has various concerns regarding the System under Consideration (SuC), also known as the System of Interest (SoI), in relation to its environment. Concerns arise throughout the life cycle of the system from the system requirements, design choices, implementation, and operations. Performance, reliability, security, privacy, distribution, openness, evolvability, modularity, cost, and regulatory compliance are a number of examples of concerns [1, 3]. Furthermore, Separation of Concerns (SoC), which is a vital design principle in SE, can be applied at different abstraction levels. At a lower level, it is interpreted as the modularity of the software system implementation, with information hiding and encapsulation in modules that have well-defined interfaces. However, at a higher level, it means describing the system architecture from the perspective of different sets of concerns (i.e., different architecture viewpoints). This is what an architecture view does. It is a work product that expresses ‘the architecture of a system from the perspective of specific system concerns’ [1]. Additionally, an architecture viewpoint is a work product that establishes ‘the conventions for the construction, interpretation, and use of architecture views to frame specific system concerns’ [1]. In other words, a view is what a stakeholder can see, whereas a viewpoint is where the stakeholder is looking from (i.e., the perspective or template that determines what they should see). A viewpoint is generic and can be stored in a library for reuse. However, a view is specific to the architecture for which it is created [3]. For instance, Table I and Figure 1 illustrate a sample architecture viewpoint and its corresponding architecture view, respectively, based on the TOGAF architecture framework. The view describes the architecture of a Cyber-Physical System (CPS) concerning the physical location of the sensors in the distributed system.
TABLE I: A sample architecture viewpoint in TOGAF [3]
@@ II Background
Figure 1: A sample architecture view in TOGAF [3] governed by the viewpoint of Table I
@@
The “4+1” View Model of Software Architecture [8], which was proposed in 1995 and updated in 2009 [9], considers the following architecture viewpoints [^1] Note that they used the term view in their work, which corresponded to the notion of viewpoint as defined by ISO/IEC/IEEE 42010:2011 [1]. for software systems: i) The logical viewpoint concentrates on the functionality of the software system. The key stakeholder group whose concerns are framed by this viewpoint is the end-user group. Moreover, the components in the views associated with this viewpoint are often classes. ii) The process viewpoint focuses on the non-functional aspects, such as performance, availability, fault tolerance, integrity, and scalability. This viewpoint is considered primarily for system integrators and system engineers. Furthermore, the components in the views associated with this viewpoint are typically tasks. iii) The development viewpoint is concerned with the software development and management aspects, such as the subsystems and modules’ organization, reuse, and portability. The main stakeholder groups here are the programmer group (i.e., software developers) as well as software managers. Also, the components in the views associated with this viewpoint are usually modules and subsystems. iv) The physical viewpoint deals with scalability, performance, and availability on the physical layer, thus considering the network communications, distribution, topology, and the mapping of software onto the hardware. The main stakeholder group here is the system engineer group. Moreover, the components in the views associated with this viewpoint are often nodes. v) The scenarios viewpoint ensures understandability and includes a small set of use case scenarios that can show how the elements of the above-mentioned viewpoints can work together. In fact, this viewpoint is an abstraction of the system requirements. It is worth mentioning that this viewpoint is in some sense redundant from the other ones. This is why the name contains ‘+1’. However, it serves the important purpose of being a driver to discover the architectural elements during the design phase and playing a validation and illustration role after the completion of the design for the test of an initial system prototype. This viewpoint is mainly intended for end-users and software developers. Lastly, the components of the views associated with this viewpoint are steps and scripts. vi) The decision viewpoint: Architects have various design choices and make a lot of design decisions. These decisions must be rigorously documented. The focus of this new viewpoint in the revised version of the original ‘4+1’ model is on design decisions.
@@
Further, the aforementioned standard [1] defined a model kind as a set of conventions for a type of modeling. An architecture viewpoint comprises one or more model kinds. Similarly, an architecture view comprises one or more architecture models. Additionally, each and every architecture viewpoint governs one or more architecture views; also, each and every model kind governs one or more architecture models. For instance, ‘data flow diagrams, class diagrams, Petri nets, balance sheets, organization charts, and state transition models’ are model kinds [1]. Finally, architecting is the ‘process of conceiving, defining, expressing, documenting, communicating, certifying proper implementation of, maintaining, and improving an architecture throughout a system life cycle’ [1]. It is typically conducted in the context of a project or an organization.
III Related Work
@@
The “4+1” View Model of Software Architecture, as explained in Section II, was an early but leading work concerned with the architecture of software systems. Other architecture frameworks, such as the Zachman Framework [7], which preceded this, or the other ones that succeeded it (see Section I), had also identified a number of stakeholders, as well as several viewpoints and views. In the following, we refer to DoDAF [4] and TOGAF [2, 3] whose latest versions were published in 2010 and 2018, respectively, as two representative examples. In particular, we highlight their data-related aspects since those are the most related ones to the target of this study.
@@
DoDAF [4] defined several viewpoints including the Data and Information Viewpoint 1 (DIV-1, called conceptual data model), DIV-2 (called logical data model), and DIV-3 (called physical data model). DIV-1 was mainly intended for non-technical stakeholders and showed the high-level data concepts and their relationships. Relationships at this level would be simple (i.e., not attributed). However, DIV-2 bridged the gap between the conceptual and physical levels by introducing the attributes and structural business process (activity) rules that formed the data structure. This viewpoint was intended for system architects and analysts. For example, an Entity-Relationship (ER) diagram or a Class diagram could be deployed as a candidate model kind for this viewpoint. Lastly, DIV-3 provided the physical schema for the data and information and was close to the actual implementation level. This could often be represented by tables, records, and keys in a relational database management system or through objects in an object-oriented data model. The physical data model (i.e., DIV-3) was intended for database engineers, software developers, and system engineers.
@@
Moreover, the TOGAF standard [2, 3] is closely related to the Zachman Framework [7] [^2] In fact, there exists a mapping [2] between the TOGAF Architecture Development Method (ADM) and the Zachman Framework.. In its 2006 version (i.e., v8.1.1), TOGAF [2] adopted the data flow viewpoint [^3] Note that they used the term view in their work, which corresponded to the notion of viewpoint as defined by ISO/IEC/IEEE 42010:2011 [1]., which was intended for database engineers, ‘concerned with the storage, retrieval, processing, archiving, and security of data’, thus ‘assuring ubiquitous access to high-quality data’ [2]. Further, the latest version of TOGAF (v9.2) from 2018 [3] promoted a set of data-related viewpoints (which they called data diagrams): i) The conceptual data viewpoint had the key purpose of depicting the relationships between the critical data entities of an enterprise. This viewpoint was intended for business stakeholders. ER diagrams or simplified UML Class diagrams could be deployed as the model kinds supporting this viewpoint. ii) The logical data viewpoint was intended for application developers and database designers and showed the logical views of the relationships between the critical data entities within an enterprise. iii) The data dissemination viewpoint illustrated the relationship between data entity, business service, and application components. It showed how the logical entities were physically realized by application components. iv) The data security viewpoint intended to depict which actor (i.e., person, organization, or system) has access to which enterprise data. This could be demonstrated in a matrix form or as a mapping. Moreover, it could show compliance with data privacy laws and other applicable regulations. v) The data migration viewpoint showed the flow of data between the source and target applications when implementing a package or packaged service-based solution, for example, if a legacy application was to be replaced. Packages might have their own data model. Thus, data transformation might be necessary. The transformation could include data quality processes, namely data cleansing, matching, merging, and consolidating data from different sources, as well as source-to-target mappings. This viewpoint could be deployed for data auditing and establishing traceability. The detail level of the supporting views could vary. vi) The data life cycle viewpoint enabled managing business data throughout their life cycle from conception until disposal. Each change in the state of data should be shown through the views supporting this viewpoint. The data were considered entities in their own rights, decoupled from business processes and activities. This allowed common data requirements to be identified.
@@
Recently, various works have pointed out the need for upgrading architecture frameworks to support ML-enabled systems. Almost around the same time that we conducted our empirical study (2021), Muccini and Vaidhyanathan [19] published a related research paper based on their experience in architecting an ML-based software system. They identified the following stakeholders with ML-related concerns: data scientists, ML developers, ethics experts, and data engineers. Moreover, they envisioned the following concerns for them: ML model accuracy, model versioning, data quality, privacy, ethics, framework, and algorithmic choice. Our empirical results confirmed their identified stakeholders and had some overlap with their stated concerns.
Additionally, Lewis et al. [20] illustrated the role of architects in ML-enabled systems and highlighted some new challenges faced by architects and other stakeholders in devising software architecture of ML-enabled software systems, namely software systems that rely on one or more ML components (including ML models) to provide their capabilities. Their study was also published in 2021 when we conducted our survey. Their work emphasized the importance of bringing the SE and data science communities together to address the concerns of both stakeholder groups holistically. They also agreed with our understanding that ML artifacts must be treated as special software artifacts, and existing principles and practices in architecture frameworks must be updated to address new concerns emerging from these artifacts.
We started from our own experience to identify the gap in the literature and formulate the research questions. However, unlike most of the related work, we conducted an empirical study that referred to domain experts and asked them for their input, positions, and preferences on the matter. Note that Lewis et al. [16] also conducted practitioner interviews as part of their work. They showed that the development and deployment of ML-enabled systems involve three different perspectives, which include data science, software engineering, and operations. If they are misaligned due to incorrect assumptions, they cause ML mismatches, which can result in failed systems. Their study was also published in 2021 when we surveyed our experts.
IV Research design and methodology